The larger of two numbers is 23. If three times the larger is 5 more than eight times the smaller, find the smaller number.
Answer by Maths68(1474)

Let
Smaller number = x
Given
Larger = y = 23
three times the larger is 5 more than eight times the smaller
8x+5 = 3y
8x = 3y - 5
8x=3(23)-5
8x=69-5
8x=64
x=64/8
x=8
Smaller number is 8